HELP what is the ratio of carboxylate anion to carboxylic acid for a carboxylate acid with...

HELP

what is the ratio of carboxylate anion to carboxylic acid for a carboxylate acid with a pKa=3.5 at pH=8?
Answer #1


using the henderson equation

pH = pKa +log[A-] / [HA]

where [HA] = weak acid concentration

[A-] = conjugate base concentration

we know pH = 8 and pKa = 3.5 substitute the values in the equation

pH = pKa +log[A-] / [HA]

8 = 3.5 +log[A-] / [HA]

log[A-] / [HA] = 4.5

then => [A-] / [HA] = 31622.77 = 3.16*104

answer => the ratio is = 31622.77 = 3.16*104